# Towards a generic absoluteness theorem for Chang models

### v2

@@ 10.1. Resurrection of derived models
Suppose $\kappa$ is a supercompact cardinal and there is a proper class of inaccessible limits of Woodin cardinals. Suppose $\lambda$ is an inaccessible limit of Woodin cardinals above $\kappa$, $\delta>\lambda$ is a Woodin cardinal and $\chi>\delta$ is an inaccessible cardinal. Let $j:V\rightarrow M$ be a $\chi$-supercompactness embedding, i.e., ${\rm cp}(j)=\kappa$, $j(\kappa)>\chi$ and $M^{\chi}\subseteq M$. Suppose $g\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic, $\mathbb{P}\in V_{\chi}[g]$ is such that $V[g]\models``\mathbb{P}$ is countably closed”, $k\subseteq\mathbb{P}$ is $V[g]$-generic and $(\pi,N,\sigma,h)\in V[g*k]$ is such that
@@
    $\sigma:N\rightarrow j(V_{\chi})$, and $j\restriction V_{\chi}=\sigma\circ\pi$, and
@@
Let $p=(f,W_{1},W_{0})\in V_{\chi}$ be a flipping system (see Definition 2.8) that flips through $\delta$, ${\sf{comp}}(p)>\lambda^{+}$ and $W_{1}$ is $(\chi,\lambda^{+},\lambda)$-saturated (see Definition 4.9). We have that if $r$ is ${<}\lambda$-generic and $\bar{\mu}\subseteq W_{1}$ is a homogeneity system in $V[r]$ then $S_{\bar{\mu}_{r}}\in\Gamma^{\infty}_{r}$. This follows from Lemma 2.11 and is also proven in [6] (e.g. see [6, Lemma 3.4.11]). Notice that because $M[g*k]$ is countably closed in $V[g*k]$, we have that $(\pi,N,\sigma,h)\in M[g*k]$.
@@
(a) $D\in N[h]$ and for some $\bar{\nu}\subseteq\pi(W_{1})$ with $\bar{\nu}\in N[h]$, $N[h]\models D=S_{\bar{\nu}}$.
Let $\zeta\in(\xi,\lambda)$ be a Woodin cardinal and let $W^{*}_{1}\subseteq W_{1}$ be $(\chi,\zeta,\xi)$-saturated and such that $\bar{\mu}\subseteq W^{*}_{1}$. Let $q=(f^{*},W^{*}_{1},W^{*}_{0})$ be a flipping system that flips through $\zeta$ and such that ${\sf{comp}}(q)>\xi$.
@@
Next we introduce genericity iterations, which are the ways models such as $N$ in Lemma 10.1 are produced. We will use Neeman’s genericity iterations as in previous sections of the paper.
@@
    for each limit ordinal $\xi<\lambda$, ${\mathcal{M}}_{\xi}$ is the direct limit of $({\mathcal{M}}_{\beta},\pi_{\beta,\beta^{\prime}}\mid\beta<\beta^{\prime}<\xi)$ where $\pi_{\beta,\beta^{\prime}}:{\mathcal{M}}_{\beta}\rightarrow{\mathcal{M}}_{\beta^{\prime}}$ is the composition of the iteration embeddings given by the iteration trees $({\mathcal{X}}_{\beta}\mid\beta<\xi)$,
  4. (GI.4)  [^44] We have that $\pi_{\beta,\beta^{\prime}}(\lambda)=\lambda$.
@@
    for each $\xi<\lambda$, ${\mathcal{X}}_{\xi}$ is a nice iteration tree of length $\omega+1$ on ${\mathcal{M}}_{\xi}$ that is below $\pi_{0,\xi}(\eta_{\xi})$, is above $\sup_{\beta<\xi}\pi_{0,\xi}(\eta_{\beta})$ and for some ${\mathcal{M}}^{{\mathcal{X}}_{\xi}}$-generic $h\subseteq\operatorname{Col}(\omega,\pi_{0,\xi+1}(\eta_{\xi}))$ with $h\in V[g]$, $a_{\xi}\in{\mathcal{M}}^{{\mathcal{X}}_{\xi}}[h]$.
@@
Our next lemma shows that the uB powerset can be correctly computed over the iterates of $V$.
@@
Suppose $\kappa$ is a supercompact cardinal and there is a proper class of inaccessible limits of Woodin cardinals. Suppose $\lambda$ is an inaccessible limits of Woodin cardinals above $\kappa$ and suppose $\delta_{0}$ is a Woodin cardinal and $\delta_{0}<\delta_{1}$ is an inaccessible cardinal such that $\delta_{0}>\lambda$. Let $j:V\rightarrow M$ be a $\delta_{1}$-supercompactness embedding, i.e., ${\rm cp}(j)=\kappa$, $j(\kappa)>\delta_{1}$ and $M^{\delta_{1}}\subseteq M$. Suppose $g\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic, $\lambda_{0}<\lambda$ is such that $\lambda_{0}=\omega_{1}^{V[g_{\lambda_{0}}]}$ and $(\pi,N,\sigma)\in V[g]$ is such that
@@
    $\pi:V_{\delta_{1}}\rightarrow N$,
@@
    $h$ appears in a forcing extension of $V[g_{\lambda_{0}}]$ by some poset $\mathbb{P}$ such that ${|\mathbb{P}|}^{V[g_{\lambda_{0}}]}=\lambda_{0}$ [^45] See [6, Lemma 3.1.5]. As was commented before, the generic $h$ is added by the poset $\mathbb{P}$ appearing in the proof of [6, Lemma 3.1.5]., and
@@
Suppose $\Gamma^{\infty}_{g_{\lambda_{0}}}=(\Gamma^{\infty})^{N[h]}$ and $X\in V_{\lambda_{1}}[g_{\lambda_{0}}]\cap V_{\lambda_{1}}^{N[h]}$. Then ${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}={\wp}_{uB}(X)^{N[h]}$.
@@
Let $\iota=\max(\iota_{X}^{N[h]},\iota_{X})$. Because $\lambda_{1}$ is an inaccessible cardinal, we have that $\iota<\lambda_{1}$. We intend to compute ${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}$ in $V[g_{\lambda_{0}}]$ after forcing with $\operatorname{Col}(\omega,\iota)$, and since ${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}$ is independent of the generic we pick for $\operatorname{Col}(\omega,\iota)$, we may just as well assume that our generic is in $V[g]$. Let then $k\in{\wp}(\operatorname{Col}(\omega,\iota))\cap V[g]$ be both $V[g_{\lambda_{0}}]$-generic and $N[h]$-generic.
Let $r\subseteq\operatorname{Col}(\omega_{1}^{V[g]},\mathbb{R}_{g})$ be $V[g]$-generic. Notice that because $M[g]$ is countably closed in $V[g]$, we have that $(\pi,N,\sigma)\in M[g]$. We now perform an $\mathbb{R}_{g}$-genericity iteration of $N$ (in $M[g*r]$) and obtain $(\pi^{\prime},N^{\prime},\sigma^{\prime})\in M[g*r]$ such that
@@
    $\pi^{\prime}:N\rightarrow N^{\prime}$ is an elementary embedding,
@@
    $\pi^{\prime}(\lambda_{1})=\lambda_{1}$ and ${\rm cp}(\pi^{\prime})>\xi$ where $\xi$ is such that $({\wp}_{uB}(X))^{N[h]}\in V_{\xi}^{N[h]}$,
@@
    every $x\in\mathbb{R}_{g}$ is in some ${<}\lambda_{1}$-generic extension of $N^{\prime}[h]$,
@@
Let now $h^{\prime}\subseteq\operatorname{Col}(\omega,{<}\lambda_{1})$ be an $N^{\prime}[h]$-generic with $k\in N^{\prime}[h*h^{\prime}]$ and $\mathbb{R}^{N^{\prime}[h*h^{\prime}]}=\mathbb{R}_{g}$. We can find such an $h^{\prime}$ in a forcing extension of $V[g*r]$ by some poset $\mathbb{P}$ such that ${|\mathbb{P}|}^{V[g]}=\lambda_{1}$ and $\mathbb{P}$ is countably closed in $V[g*r]$ (see Remark 6.7). It then follows from Lemma 10.1 that $(\Gamma^{\infty})^{N^{\prime}[h*h^{\prime}]}=\Gamma^{\infty}_{g}$. Let $j_{h,h^{\prime}}^{N^{\prime}[h*h^{\prime}]}\colon L(\Gamma_{h}^{\infty},\mathbb{R}_{h})^{N^{\prime}[h]}\rightarrow L(\Gamma^{\infty}_{h*h^{\prime}},\mathbb{R}_{h*h^{\prime}})^{N^{\prime}[h*h^{\prime}]}$ be the canonical embedding [^46] There are many such embeddings. By “the canonical” embedding, we mean the embedding that moves indiscernibles to indiscernibles in the order preserving way. in $N^{\prime}[h*h^{\prime}]$ with the property that $j_{h,h^{\prime}}(A)=A_{h^{\prime}}$ for $A\in\Gamma_{h}^{\infty}$. Moreover, let $j_{g_{\lambda_{0}},g}\colon L(\Gamma_{g_{\lambda_{0}}}^{\infty},\mathbb{R}_{g_{\lambda_{0}}})\rightarrow L(\Gamma^{\infty}_{g_{\lambda_{0}}*g},\mathbb{R}_{g_{\lambda_{0}}*g})$ be the canonical embedding with the property that $j_{g_{\lambda_{0}},g}(A)=A_{g}$ for $A\in\Gamma_{g_{\lambda_{0}}}^{\infty}$. Notice that
@@
Here we prove a derived model theorem for the uB powerset. The proof is very much like the proof of the $\sf{Derived\ Model\ Theorem}$ (see [26]). Again, we do not use the terminology of Theorem 1.19. Recall that $\mathcal{A}^{\infty}_{h}=({\wp}_{uB}(\Gamma^{\infty}))^{V[h]}$.
@@
Suppose $\kappa$ is a supercompact cardinal, there is a proper class of inaccessible limits of Woodin cardinals and $\lambda$ is an inaccessibe limit of Woodin cardinals above $\kappa$. Suppose $h\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic. Then $L(\mathcal{A}^{\infty}_{h})\models{\sf{AD^{+}}}$.
@@
The definition of $(\nu,\eta,\eta^{\prime},\bar{\mu}_{0},\bar{\mu}_{1})$.
Let $\nu<\lambda$ be such that there are $\bar{\mu}^{0},\bar{\mu}^{1}\in V[h_{\nu}]$ such that
@@
    all models of ${\mathcal{T}}$ are well-founded.
It follows from Windszus’ Lemma (see [26, Lemma 1.1]) that in $M[h_{\eta^{\prime}}]$, $\Sigma$ is $j(\nu)$-homogeneously Suslin, and since $j(\nu)>\lambda_{0}$, we have that $\Sigma_{h}\in\Gamma^{\infty}_{h}$. The following is our main claim. It implies $C\in\Gamma^{\infty}_{h}$ as $C$ is defined by real quantification from $\Sigma_{h}\in\Gamma^{\infty}_{h}$ and will therefore finish the proof of Theorem 10.4.
@@
Because ${\mathcal{T}}\in\Sigma_{h}$, we have (in $V[h]$) an embedding $\sigma_{0}^{\prime}:{\mathcal{M}}^{\mathcal{T}}\rightarrow j(V_{\eta^{\prime}})$ such that $j\restriction V_{\eta^{\prime}}=\sigma_{0}^{\prime}\circ\pi^{\mathcal{T}}$. Let now ${\mathcal{U}}$ be the result of applying ${\mathcal{T}}$ to $V_{\delta_{1}}$ [^51] This means that we apply the extenders of ${\mathcal{T}}$ to $V_{\delta_{1}}$ rather than $V_{\eta^{\prime}}$.. We then have $\sigma_{0}:{\mathcal{M}}^{\mathcal{U}}\rightarrow j(V_{\delta_{1}})$ such that $\sigma_{0}^{\prime}=\sigma_{0}\restriction{\mathcal{M}}^{\mathcal{T}}$ and $j\restriction V_{\delta_{1}}=\sigma_{0}\circ\pi^{\mathcal{U}}$. Notice that ${\mathcal{M}}^{\mathcal{T}}=(V_{\pi^{\mathcal{U}}(\eta^{\prime})})^{{\mathcal{M}}^{\mathcal{U}}}$. Set $N_{0}={\mathcal{M}}^{\mathcal{U}}$ and $\pi_{0}=\pi^{\mathcal{U}}$. We then have that $(\pi_{0},N_{0},\sigma_{0})$ is as in the hypothesis of Definition 10.2, and so we let $(\pi,N,\sigma)$ be an $\mathbb{R}_{h}$-genericity iteration of $(\pi_{0},N_{0},\sigma_{0})$ with the property that ${\rm cp}(\pi)>\pi_{0}(\eta^{\prime})$. Let $m_{1}\subseteq\operatorname{Col}(\omega,{<}\lambda)$ be $N$-generic given by [6, Lemma 3.1.5] such that
@@
We use the notation of Theorem 1.19. We have that $\sf{Sealing}$ holds (see Theorem 1.9). Suppose $\lambda_{0}<\lambda_{1}<\chi$ are inaccessible cardinals above $\kappa$ that are limits of Woodin cardinals. Let $W\in V_{\chi}$ be infinitely tamed $(\chi,\lambda_{1}^{+},\lambda_{1})$-saturated set of measures, and let $X\prec V_{\chi+1}$ be such that $V_{\kappa+1}\cup\{\lambda_{0},\lambda_{1},W\}\subseteq X$ and $V\models{|X|}={|V_{\kappa+1}|}$. Let $j:V\rightarrow N$ be a $\chi$-supercompactness embedding and let $h\subseteq\operatorname{Col}(\omega,{<}\lambda_{0})$ be $V[g]$-generic. Then it follows from Theorem 10.4 that $L(\mathcal{A}^{\infty}_{g*h})\models{\sf{AD^{+}}}$. Let $k\subseteq\operatorname{Col}(\omega,\Gamma^{\infty}_{g*h})$ be $V[g*h]$-generic. We first claim that the following holds:
@@
be the underlying sequence of ${\sf{DM}}$, $M_{\omega}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega}:M_{\omega}\rightarrow N_{j(\chi)}$ be the canonical realizable embedding. As in the proof of Theorem 6.5, we can now find $M_{\omega}$-generic $m\subseteq\operatorname{Col}(\omega,{<}\lambda_{0})$ [^53] We have $\lambda_{0}=\kappa_{\omega}=\omega_{1}^{V[g*h]}$. such that $(\Gamma^{\infty})^{M_{\omega}[m]}=\Gamma^{\infty}_{g*h}$. We then claim that
@@
Clearly, given Theorem 10.4, (b) implies (a). To show (b) we apply the extenders $E_{n}$ to $V$. More precisely, let $(K_{i}:i\leq\omega)$ be obtained by setting $K_{0}=V_{\chi}$ and letting $K_{i+1}=Ult(K_{i},E_{i})$ and $K_{\omega}$ be the direct limit of $(K_{i},n_{i,i^{\prime}}:i<i^{\prime}<\omega)$ where $n_{i,i^{\prime}}:K_{i}\rightarrow K_{i^{\prime}}$ is the composition of ultrapower embeddings. Let $n_{i,\omega}$ be the direct limit embedding. The construction of the sequence $(K_{i}:i\leq\omega)$ is possible because $V_{\kappa+1}\subseteq M_{0}$, which by induction implies that for each $i\leq\omega$,
@@
be the underlying sequence of ${\sf{DM}}$, $M_{\omega\cdot 2}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega\cdot 2}:M_{\omega\cdot 2}\rightarrow N_{j(\chi)}$ be the canonical realizable embedding.
@@
The proof of Theorem 1.30 is an application of Theorem 6.6 and [30, Theorem 0.1]. We start by proving the first part of the Sealing Theorem in the sense of Theorem 1.30. This means that we start by establishing Clauses 1 and 2 of Definition 1.24 for the model $L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}^{\infty}]$. The proof of Clause 3 of Definition 1.24 follows the very same path as the proof of Clause 3 in Theorem 1.9 (see Section 7) and in Theorem 1.19 (see Subsection 1.19). Therefore, we will only indicate how to prove it.
To start with notice that $j(\kappa)$ satisfies the hypothesis of [30, Theorem 0.1] (in $V^{*}$, with $j(\kappa)=\delta_{0}$). Therefore, the following holds by the proof of [30, Theorem 0.1].

### v3

@@ 10.1. Resurrection of derived models
Suppose $\kappa$ is a supercompact cardinal and there is a proper class of inaccessible limits of Woodin cardinals. Suppose $\lambda$ is an inaccessible limit of Woodin cardinals above $\kappa$, $\delta>\lambda$ is a Woodin cardinal and $\chi>\delta$ is an inaccessible cardinal. Let $j:V\rightarrow M$ be a $\chi$-supercompactness embedding, i.e., ${\rm cp}(j)=\kappa$, $j(\kappa)>\chi$ and $M^{\chi}\subseteq M$. Suppose $g\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic, $\mathbb{P}\in V_{\chi}[g]$ is such that $V[g]\models``\mathbb{P}$ is countably closed”, $k\subseteq\mathbb{P}$ is $V[g]$-generic and $(\pi,N,\sigma,h)\in V[g*k]$ are such that
@@
    $\sigma:N\rightarrow j(V_{\chi})$, $j\restriction V_{\chi}=\sigma\circ\pi$, and
@@
Let $p=(f,W_{1},W_{0})\in V_{\chi}$ be a flipping system (see Definition 2.8) that flips through $\delta$ such that ${\sf{comp}}(p)>\lambda^{+}$ and $W_{1}$ is $(\chi,\lambda^{+},\lambda)$-saturated (see Definition 4.9). We have that if $r$ is ${<}\lambda$-generic and $\bar{\mu}\subseteq W_{1}$ is a homogeneity system in $V[r]$ then $S_{\bar{\mu}_{r}}\in\Gamma^{\infty}_{r}$. This follows from Lemma 2.11 and is also proven in [7] (e.g. see [7, Lemma 3.4.11]). Notice that because $M[g*k]$ is $\chi$-closed in $V[g*k]$, we have that $(\pi,N,\sigma,h)\in M[g*k]$.
@@
(a) $D\in N[h]$ and for some homogeneity system $\bar{\nu}\subseteq\pi(W_{1})$ with $\bar{\nu}\in N[h]$, $N[h]\models D=S_{\bar{\nu}}$.
Let $\zeta\in(\xi,\lambda)$ be a Woodin cardinal and let $W^{*}_{1}\subseteq W_{1}$ be $(\chi,\zeta,\xi)$-saturated and such that $\bar{\mu}\subseteq W^{*}_{1}$. Let $q=(f^{*},W^{*}_{1},W^{*}_{0})$ be a flipping system that flips through $\zeta$ with ${\sf{comp}}(q)>\xi$.
@@
Next we introduce genericity iterations, which are the ways models such as $N$ in Lemma 10.1 are produced. We will use Neeman’s genericity iterations (see Theorem 2.5) as in previous sections of the paper.
@@
    for each limit ordinal $\xi<\lambda$, ${\mathcal{M}}_{\xi}$ is the direct limit of $({\mathcal{M}}_{\beta},\pi_{\beta,\beta^{\prime}}\mid\beta<\beta^{\prime}<\xi)$ where $\pi_{\beta,\beta^{\prime}}:{\mathcal{M}}_{\beta}\rightarrow{\mathcal{M}}_{\beta^{\prime}}$ is the composition of the iteration embeddings given by the iteration trees $({\mathcal{X}}_{\beta}\mid\beta<\xi)$, [^47] We have that $\pi_{\beta,\beta^{\prime}}(\lambda)=\lambda$.
  4. (GI.4)
@@
    for each $\xi<\lambda$, ${\mathcal{X}}_{\xi}$ is a nice iteration tree of length $\omega+1$ on ${\mathcal{M}}_{\xi}$ that is below $\pi_{0,\xi}(\eta_{\xi})$, above $\sup_{\beta<\xi}\pi_{0,\xi}(\eta_{\beta})$ and for some ${\mathcal{M}}^{{\mathcal{X}}_{\xi}}$-generic $h\subseteq\operatorname{Col}(\omega,\pi_{0,\xi+1}(\eta_{\xi}))$ with $h\in V[g]$, $a_{\xi}\in{\mathcal{M}}^{{\mathcal{X}}_{\xi}}[h]$.
@@
Our next lemma shows that the uB-powerset can be correctly computed over the iterates of $V$.
@@
Suppose $\kappa$ is a supercompact cardinal and there is a proper class of inaccessible limits of Woodin cardinals. Suppose $\lambda$ is an inaccessible limit of Woodin cardinals above $\kappa$ and suppose $\delta_{0}$ is a Woodin cardinal and $\delta_{0}<\delta_{1}$ is an inaccessible cardinal such that $\delta_{0}>\lambda$. Let $j:V\rightarrow M$ be a $\delta_{1}$-supercompactness embedding, i.e., ${\rm cp}(j)=\kappa$, $j(\kappa)>\delta_{1}$ and $M^{\delta_{1}}\subseteq M$. Suppose $g\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic, $\lambda_{0}<\lambda$ is such that $\lambda_{0}=\omega_{1}^{V[g_{\lambda_{0}}]}$ and $(\pi,N,\sigma)\in V[g]$ is such that
@@
    $\pi:V_{\delta_{1}}\rightarrow N$ and $N\subseteq V_{\delta_{1}}[g]$,
@@
    $h$ appears in a forcing extension of $V[g_{\lambda_{0}}]$ by some poset $\mathbb{P}$ such that ${|\mathbb{P}|}^{V[g_{\lambda_{0}}]}=\lambda_{0}$ [^48] See [7, Lemma 3.1.5]. As was commented before, the generic $h$ is added by the poset $\mathbb{P}$ appearing in the proof of [7, Lemma 3.1.5]., and
@@
Suppose $\Gamma^{\infty}_{g_{\lambda_{0}}}=(\Gamma^{\infty})^{N[h]}$ and $X\in V_{\lambda}[g_{\lambda_{0}}]\cap V_{\lambda}^{N[h]}$. Then
${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}={\wp}_{uB}(X)^{N[h]}.$
@@
Let $\iota=\max(\iota_{X}^{N[h]},\iota_{X})$. Because $\lambda$ is an inaccessible cardinal, we have that $\iota<\lambda$. We intend to compute ${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}$ in $V[g_{\lambda_{0}}]$ after forcing with $\operatorname{Col}(\omega,\iota)$, and since ${\wp}_{uB}(X)^{V[g_{\lambda_{0}}]}$ is independent of the generic we pick for $\operatorname{Col}(\omega,\iota)$, we may just as well assume that our generic is in $V[g]$. Let then $k\in{\wp}(\operatorname{Col}(\omega,\iota))\cap V[g]$ be both $V[g_{\lambda_{0}}]$-generic and $N[h]$-generic.
Let $r\subseteq\operatorname{Col}(\omega_{1}^{V[g]},\mathbb{R}_{g})$ be $V[g]$-generic. Notice that because $M[g]$ is $\delta_{1}$-closed in $V[g]$, we have that $(\pi,N,\sigma)\in M[g]$. We now perform an $\mathbb{R}_{g}$-genericity iteration of $N$ (in $M[g*r]$) and obtain $(\pi^{\prime},N^{\prime},\sigma^{\prime})\in M[g*r]$ such that
@@
    $\pi^{\prime}:N\rightarrow N^{\prime}$ is an elementary embedding and $N^{\prime}\subseteq V_{\delta_{1}}[g*r]$,
@@
    $\pi^{\prime}(\lambda)=\lambda$ and ${\rm cp}(\pi^{\prime})>\xi$ where $\xi$ is such that $({\wp}_{uB}(X))^{N[h]}\in V_{\xi}^{N[h]}$,
@@
    every $x\in\mathbb{R}_{g}$ is in some ${<}\lambda$-generic extension of $N^{\prime}[h]$,
@@
Let now $h^{\prime}\subseteq\operatorname{Col}(\omega,{<}\lambda)$ be $N^{\prime}[h]$-generic with $k\in N^{\prime}[h*h^{\prime}]$ and $\mathbb{R}^{N^{\prime}[h*h^{\prime}]}=\mathbb{R}_{g}$. We can find such an $h^{\prime}$ in a forcing extension of $V[g*r]$ by some poset $\mathbb{P}$ such that ${|\mathbb{P}|}^{V[g]}=\lambda$ and $\mathbb{P}$ is countably closed in $V[g*r]$ (see Remark 6.7). It then follows from Lemma 10.1 that $(\Gamma^{\infty})^{N^{\prime}[h*h^{\prime}]}=\Gamma^{\infty}_{g}$. Let $j_{h,h^{\prime}}^{N^{\prime}[h*h^{\prime}]}\colon L(\Gamma_{h}^{\infty},\mathbb{R}_{h})^{N^{\prime}[h]}\rightarrow L(\Gamma^{\infty}_{h*h^{\prime}},\mathbb{R}_{h*h^{\prime}})^{N^{\prime}[h*h^{\prime}]}$ be the canonical embedding [^49] There are many such embeddings. By “the canonical” embedding, we mean the embedding that moves indiscernibles to indiscernibles in the order preserving way. in $N^{\prime}[h*h^{\prime}]$ with the property that $j_{h,h^{\prime}}(A)=A_{h^{\prime}}$ for $A\in\Gamma_{h}^{\infty}$. Moreover, let $j_{g_{\lambda_{0}},g}\colon L(\Gamma_{g_{\lambda_{0}}}^{\infty},\mathbb{R}_{g_{\lambda_{0}}})\rightarrow L(\Gamma^{\infty}_{g_{\lambda_{0}}*g},\mathbb{R}_{g_{\lambda_{0}}*g})$ be the canonical embedding with the property that $j_{g_{\lambda_{0}},g}(A)=A_{g}$ for $A\in\Gamma_{g_{\lambda_{0}}}^{\infty}$. Notice that
@@
Here we prove a derived model theorem for the uB-powerset. The proof is very much like the proof of the $\sf{Derived\ Model\ Theorem}$ (see [26]). Again, we do not use the terminology introduced in the previous section. Recall that $\mathcal{A}^{\infty}_{h}=({\wp}_{uB}(\Gamma^{\infty}))^{V[h]}$.
@@
Suppose $\kappa$ is a supercompact cardinal, there is a proper class of inaccessible limits of Woodin cardinals and $\lambda$ is an inaccessible limit of Woodin cardinals above $\kappa$. Suppose $h\subseteq\operatorname{Col}(\omega,{<}\lambda)$ is $V$-generic. Then $L(\mathcal{A}^{\infty}_{h})\models{\sf{AD^{+}}}$.
@@
The definition of $(\nu,\eta,\eta^{\prime},\bar{\mu}^{0},\bar{\mu}^{1})$.
Let $\nu<\lambda$ be such that there are homogeneity systems $\bar{\mu}^{0},\bar{\mu}^{1}\in V[h_{\nu}]$ such that
@@
    all models of $j{\mathcal{T}}$, the copy of ${\mathcal{T}}$ on $M$, are well-founded.
It follows from Windszus’ Lemma (see Lemma 4.1 or [26, Lemma 1.1]) that in $M[h_{\eta^{\prime}}]$, $\Sigma$ is $j(\nu)$-homogeneously Suslin, and since $j(\nu)>\lambda_{0}$, we have that $\Sigma_{h}\in\Gamma^{\infty}_{h}$. The following is our main claim. It implies $C\in\Gamma^{\infty}_{h}$ as $C$ is defined by real quantification from $\Sigma_{h}\in\Gamma^{\infty}_{h}$ and will therefore finish the proof of Theorem 10.4.
@@
Because ${\mathcal{T}}\in\Sigma_{h}$, we have (in $V[h]$) an embedding $\sigma_{0}^{\prime}:{\mathcal{M}}^{\mathcal{T}}\rightarrow j(V_{\eta^{\prime}})$ such that $j\restriction V_{\eta^{\prime}}=\sigma_{0}^{\prime}\circ\pi^{\mathcal{T}}$. Let now ${\mathcal{U}}$ be the result of applying ${\mathcal{T}}$ to $V_{\delta_{1}}$ [^55] This means that we apply the extenders of ${\mathcal{T}}$ to $V_{\delta_{1}}$ rather than $V_{\eta^{\prime}}$.. We then have $\sigma_{0}:{\mathcal{M}}^{\mathcal{U}}\rightarrow j(V_{\delta_{1}})$ such that $\sigma_{0}^{\prime}=\sigma_{0}\restriction{\mathcal{M}}^{\mathcal{T}}$ and $j\restriction V_{\delta_{1}}=\sigma_{0}\circ\pi^{\mathcal{U}}$. Notice that ${\mathcal{M}}^{\mathcal{T}}=(V_{\pi^{\mathcal{U}}(\eta^{\prime})})^{{\mathcal{M}}^{\mathcal{U}}}$. Set $N_{0}={\mathcal{M}}^{\mathcal{U}}$ and $\pi_{0}=\pi^{\mathcal{U}}$. We then have that $(\pi_{0},N_{0},\sigma_{0})$ is as in the hypothesis of Definition 10.2, and so we let $(\pi,N,\sigma)$ be an $\mathbb{R}_{h}$-genericity iteration of $(\pi_{0},N_{0},\sigma_{0})$ with the property that ${\rm cp}(\pi)>\pi_{0}(\eta^{\prime})$. Let $m_{1}\subseteq\operatorname{Col}(\omega,{<}\lambda)$ be $N$-generic given by [7, Lemma 3.1.5] such that
@@
We use the notation of Theorem 1.19. We have that $\sf{Sealing}$ holds (see Theorem 1.9). Suppose $\lambda_{0}<\lambda_{1}<\chi$ are inaccessible cardinals above $\kappa$ that are limits of Woodin cardinals. Let $W\in V_{\chi}$ be an infinitely tamed $(\chi,\lambda_{1}^{+},\lambda_{1})$-saturated set of measures, and let $X\prec V_{\chi+1}$ be such that $V_{\kappa+1}\cup\{\lambda_{0},\lambda_{1},W\}\subseteq X$ and $V\models{|X|}={|V_{\kappa+1}|}$. Let $j:V\rightarrow N$ be a $\chi$-supercompactness embedding and let $h\subseteq\operatorname{Col}(\omega,{<}\lambda_{0})$ be $V[g]$-generic. Then it follows from Theorem 10.4 that $L(\mathcal{A}^{\infty}_{g*h})\models{\sf{AD^{+}}}$. Let $k\subseteq\operatorname{Col}(\omega,\Gamma^{\infty}_{g*h})$ be $V[g*h]$-generic. We first claim that the following holds:
@@
be the underlying sequence of $D$, $M_{\omega}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega}:M_{\omega}\rightarrow N_{j(\chi)}$ be the canonical realizability embedding. As in the proof of Theorem 6.5, we can now find an $M_{\omega}$-generic $m\subseteq\operatorname{Col}(\omega,{<}\lambda_{0})$ [^57] We have $\lambda_{0}=\kappa_{\omega}=\omega_{1}^{V[g*h]}$. such that $(\Gamma^{\infty})^{M_{\omega}[m]}=\Gamma^{\infty}_{g*h}$. We then claim that
@@
Clearly, given Theorem 10.4, (b) implies (a). To show (b) we apply the extenders $E_{n}$ of the $\sf{DM}$-sequence $D$ to $V$. More precisely, let $(K_{i}:i\leq\omega)$ be obtained by setting $K_{0}=V_{\chi}$ and letting $K_{i+1}=Ult(K_{i},E_{i})$ and $K_{\omega}$ be the direct limit of $(K_{i},n_{i,i^{\prime}}:i<i^{\prime}<\omega)$ where $n_{i,i^{\prime}}:K_{i}\rightarrow K_{i^{\prime}}$ is the composition of ultrapower embeddings. Let $n_{i,\omega}$ be the direct limit embedding. The construction of the sequence $(K_{i}:i\leq\omega)$ is possible because $V_{\kappa+1}\subseteq M_{0}$, which by induction implies that for each $i\leq\omega$,
@@
be the underlying sequence of $D$, $M_{\omega\cdot 2}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega\cdot 2}:M_{\omega\cdot 2}\rightarrow N_{j(\chi)}$ be the canonical realizability embedding.
@@
The proof of Theorem 1.30 is an application of Theorem 6.6 and [30, Theorem 0.1]. We start by proving the first part of the Sealing Theorem in the sense of Theorem 1.30. This means that we start by establishing Clauses (1) and (2) of Definition 1.24 for the model $L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}^{\infty}]$. The proof of Clause (3) of Definition 1.24 follows the very same path as the proof of the similar clause in Theorem 1.9 (see Section 7) and in Theorem 1.19 (see Subsection 10.3). Therefore, we will only indicate how to prove it.
To start, notice that $j(\kappa)$ satisfies the hypothesis of [30, Theorem 0.1] (in $V^{*}$, with $j(\kappa)=\delta_{0}$). Therefore, the following holds by the proof of [30, Theorem 0.1].